OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 4) and is parallel to y = -3x + 6. y = one thirdx + 14 thirds y = 3x + 10 y = -3x - 2 y = -one thirdx + 10 thirds

OpenStudy (anonymous):

I would appreciate your help @sushant20902179

OpenStudy (anonymous):

the line passing through point (-2,4) is parallel to y=-3x+6...it means the slope is same..

OpenStudy (anonymous):

OK, so in order to get the answer I would have to find the perpendicular form?

OpenStudy (anonymous):

y = mx +c ; m is the slope comparing y = mx +c with the equation with y = -3x + 6 ....from this we get slope i.e m=-3

OpenStudy (anonymous):

Yeah, I know that... I just don't know how to incorporate it with the question asked...

OpenStudy (anonymous):

use the point slope form

OpenStudy (anonymous):

y= mx + c ?

OpenStudy (anonymous):

(y-y1)/(x-x1) = m as we have given the point (-2,4) ; x1 = -2 and y1 = 4

OpenStudy (anonymous):

y = mx +c is the equation of the line..

OpenStudy (anonymous):

Oh man, I am so confused... This is probably the only thing that I suck in with school....

OpenStudy (anonymous):

(y-4)/(x+2) = -3 after solving this you will get y = -3x - 2

OpenStudy (anonymous):

at which point u r getting confused...

OpenStudy (anonymous):

Ok, so I would have to complete the whole slope thingy, to get the answer? Where did you get the 4 from?

OpenStudy (anonymous):

in question u have given the points (-2,4) ....from this i got the 4

OpenStudy (anonymous):

OH... Never mind, I get it!! Thank you :) PS: I gave you a medal..

OpenStudy (anonymous):

thank you for the medal...